OpenStudy (anonymous):

I need help in simplifying (3x^2 +3x -36) / (x^2-x-6) and how do you get the restrictions? I've gotten the wrong answer 3 times already :/

Parth (parthkohli):

Restriction is when the denominator is 0.

Parth (parthkohli):

So, \( \color{Black}{\Rightarrow x^2 - x - 6 \ne 0 }\) Just solve it like a quadratic equation :)

OpenStudy (anonymous):

wait huh?

Parth (parthkohli):

Do you know how to solve quadratic equations?

OpenStudy (anonymous):

i think so..

OpenStudy (anonymous):

its with that formula right?

Parth (parthkohli):

Yeah, many ways to solve.

Parth (parthkohli):

If you know the formula, then you can use that :)

OpenStudy (anonymous):

but how would that give you a restriction like what is that

Parth (parthkohli):

Well, the denominator must equal 0 right?

Parth (parthkohli):

to get the restriction, the denominator must equal zero.

OpenStudy (anonymous):

and after it equals zero...

OpenStudy (anonymous):

could you show me an example or something?

Parth (parthkohli):

Ok

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large{4 \over x} }\) x can't be 0 here.

OpenStudy (anonymous):

so the restriction cant be 4?

Parth (parthkohli):

Restriction on any value is when the denominator is 0.

Parth (parthkohli):

To find the restriction, you must equate the denominator with 0.

OpenStudy (anonymous):

maybe im just stupid but i dont get that at all, im sorry, thank you for trying though

Parth (parthkohli):

lol no, it's my duty to explain

Parth (parthkohli):

Anything over 0 is undefined.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {y \over x - 4} }\) 1) equate the denominator with 0. \( \color{Black}{\Rightarrow x - 4 = 0 }\) 2) solve for x. \( \color{Black}{\Rightarrow x = 4 }\)

Parth (parthkohli):

So, 4 is the restriction. This was an example tho.

OpenStudy (anonymous):

could you work out the problem above to show me ? and like explain as you go? cause the one above seems alot more complicated thant he examples

Parth (parthkohli):

Okay. As I told you, equate the denominator with 0. \( \color{Black}{\Rightarrow x^2 - x - 6 = 0}\)

Parth (parthkohli):

Use the formula now. \( \color{Black}{\Rightarrow x = \Large {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

Parth (parthkohli):

a = 1 b = 1 c = -6

OpenStudy (anonymous):

but dont you have to simplify (3x^2 +3x -36) / (x^2-x-6) before you can even get the restriction?

Parth (parthkohli):

No, you don't have to :)

Parth (parthkohli):

Well you can but it'd make it longer

OpenStudy (anonymous):

but what if the problem tells you to :(

Parth (parthkohli):

Then simplify it....

Parth (parthkohli):

You can simplify by factoring.

OpenStudy (anonymous):

thats part of what i dont get.;///;.......

Parth (parthkohli):

Factor the numerator and denominator first.

OpenStudy (anonymous):

-36 doesnt factor to equal +3 though for the numerator

Parth (parthkohli):

3 is common in the numerator :D

OpenStudy (anonymous):

so then how would you right that

OpenStudy (anonymous):

write**

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3(x^2 + x - 12) }\)

Parth (parthkohli):

^ you can factor this

OpenStudy (anonymous):

ohhhh ya i forgot about that way

OpenStudy (anonymous):

wait you can factor what

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3(x^2 -3x + 4x -12) }\) :D

OpenStudy (anonymous):

oh okok

OpenStudy (anonymous):

so 3((x-3)(x+4)-12) or something i forget

Parth (parthkohli):

Yeah... \(3(x + 3)(x - 4)\)

Parth (parthkohli):

And now factor the denominator: \( \color{Black}{\Rightarrow x^2 - 3x + 2x - 6}\)

OpenStudy (anonymous):

(x-3)(x+2)

Parth (parthkohli):

yep

Parth (parthkohli):

so that's it... simplified.

Parth (parthkohli):

Now as you know factoring, no need to use the formula

OpenStudy (anonymous):

but you still hae to divide the two right

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - 3)(x + 2) = 0 }\) Can you solve now? this will get the restrictions

OpenStudy (anonymous):

are the restrictions the roots of like a graph

Parth (parthkohli):

yes.

OpenStudy (anonymous):

wowwwwwww

OpenStudy (anonymous):

hahahahah i am so sorry

OpenStudy (anonymous):

but wait 3(x+3)(x+4) / (x-3)(x+2)

Parth (parthkohli):

Yeah, but still this can't simplify

OpenStudy (anonymous):

so then how would you even find the roots??

Parth (parthkohli):

Use the zero product rule

OpenStudy (anonymous):

the what

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x -3) = 0 }\) \( \color{Black}{\Rightarrow (x + 4) = 0 }\) This will get the roots that'd be the restrictions

OpenStudy (anonymous):

does the problem have to simplify to look like that for you to be able to do that?

Parth (parthkohli):

Hmm?

OpenStudy (anonymous):

like in order to get the roots the problem usually has to simplify to look like ⇒(x−3)=0 ⇒(x+4)=0 that to be able to equal it to zero

Parth (parthkohli):

You have to solve these equations to get the solutions.

OpenStudy (anonymous):

well ya which would be +3 and -4 but

Parth (parthkohli):

Yeah, those ARE the restrictions to x :)

OpenStudy (anonymous):

if you cant simplify 3(x+3)(x-4) / (x-3)(x+2) anymore than what makes the (x-3)and (x+4) part so special, wouldnt the other parts like (x+2) =0 to get -2 as a restriction too?

Parth (parthkohli):

No! No!

Parth (parthkohli):

The fraction is undefined when the denominator is 0, not the numerator.

OpenStudy (anonymous):

but the denominator isnt zero.......

Parth (parthkohli):

The denominator cannot equal 0.

OpenStudy (anonymous):

ohh ever?

Parth (parthkohli):

It can never equal zero lol zero is a bad number

Parth (parthkohli):

the numerator can be zero on the other hand

OpenStudy (anonymous):

oh ya cause then the top would equal zero too

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

the numerator is still 3(x+3) (x-4) though.... what happened to thee three

Parth (parthkohli):

if you divide zero by something, it's zero. if you divide something by zero, YOU JUST CANT

OpenStudy (anonymous):

hello?

Parth (parthkohli):

hi?

OpenStudy (anonymous):

the numerator is still 3(x+3) (x-4) though.... what happened to the three

OpenStudy (anonymous):

i get that the restrictions came from (x+3)(x-4) but in the numerator there is a three in from of the two

OpenStudy (anonymous):

does that just dissapear?

Parth (parthkohli):

Hmm?

OpenStudy (anonymous):

i still dont get it

Parth (parthkohli):

Lol jo wil explain

OpenStudy (callisto):

If you don't mind, can I start from the beginning?

OpenStudy (anonymous):

so the answer is 3(x+3)(x-4) / (x-3)(x+2) with restrictions of -3 and +4 as long as its right you guys dont have to worry about explaning it haha

OpenStudy (anonymous):

actually thatd be awesome as long as you dont mind.....

Parth (parthkohli):

GIVE JO THE MEDALS

OpenStudy (callisto):

Nope :| Do you mind re-post the question? Lag here :|

Parth (parthkohli):

Then I deserve the medals :P

OpenStudy (callisto):

Take back the medals given to me!!!

Parth (parthkohli):

Done